General Cereal common stock dividends have been growing at an annual rate of 8 percent per year over the past 10 years. Current dividends are $1.4 per share. What is the current value of a share of this stock to an investor who requires a 11 percent rate of return if the following conditions exist? Round your answers to the nearest cent.

Dividends are expected to continue growing at the historic rate for the foreseeable future.

The dividend growth rate is expected to increase to 10 percent per year.

The dividend growth rate is expected to decrease to 7 percent per year.
Answer #1

Value of share = Expected Dividend/(Required Return – growth rate)

Dividends are expected to continue growing at the historic rate for the foreseeable future.

Value = 1.4(1+8%)/(11%-8%)

= $50.4

The dividend growth rate is expected to increase to 10 percent per year.

Value = 1.4(1.1)/(11%-10%)

= $154

The dividend growth rate is expected to decrease to 7 percent per year.

Value = 1.4(1.07)/(11%-7%)

= $37.45